maria is now 3 years older than her sister, Ann. Four years ago, Maria's age was one-half Anna's age.How old is each girl now?
Answer by ValorousDawn(53)

Let Maria's age be m, Ann's age be a.
Maria is now 3 years older than her sister, Ann. This gives us m=a+3.

Four years ago, Maria's age was one-half Anna's age. This gives us (m-4)/2=a-4. Divided by two for the one half part. Both m and a have 4 subtracted because both ages are from 4 years ago.

We can invoke substitution for this problem. Since m=a+3, we get a+3-4/2=a-4.
a-1=2a-8
a=7
If Ann's age is 7, and Maria is three years older, then Maria must be 10.